Each plate of a parallel-plate capacitor is a square of side 0.0313 m and the plates are separated by 0.423
Answer #1

Area

A=S2=0.03132=9.8*10-4 m2

Capacitance

C=eoA/d=(8.85*10-12)(9.8*10-4)/(0.423*10-3)

C=2.05*10-11 F

Energy stored in capacitor

E=(1/2)CV2

=>V=sqrt[2*E/C]=sqrt[2*8.83*10-9/2.05*10-11]

V=29.35 volts

Electric field

E=V/d=29.35/(0.423*10-3)

E=6.94*104 N/C
Answer #2

Since C=e0A/d=8.85x10-12 x 0.0313x0.0313/0.000423=20.5x10-12F

also

U=1/2CV2

V=sqrt(2U/C)=sqrt( 2 x 8.83x10-9/20.5x10-12)=29.35V

E=V/d=29.35/0.000423=69387N/C
Answer #3

my answer is 69576 N/C
Answer #4

true answer is 6.9 * 104 N/C